# STaR: Seamless Spatial-Temporal Aware Motion Retargeting with Penetration and Consistency Constraints

### v1

@@ 4.3 Ablation Study
Table 2: Ablation Study. Baseline is a model which retargets motion frame-by-frame, training with $\mathcal{L}_{sdf}$ from [20]. ST is the spatio-temporal model, and DSR stands for dense shape representation. $\mathcal{L}_{lp}$ is the limb penetration constraint, and $\mathcal{L}_{tc}$ is the temporal consistency constraint. $\mathcal{L}_{t}$ is a basic motion smoothness loss. In the second part of the table, we sample different number of vertices for limb penetration constraint.
@@
Baseline w/ $\mathcal{L}_{sdf}$ [20] 0.0479 0.0321 8.74 66.35
ST w/ $\mathcal{L}_{sdf}$ [20] 0.0371 0.0179 9.45 7.87
ST w/ $\mathcal{L}_{lp}$ 0.0361 0.0167 8.20 35.74
ST w/ $\mathcal{L}_{lp}$ w/ $\mathcal{L}_{t}$ 0.0361 0.0167 8.40 35.74
ST w/ DSR w/ $\mathcal{L}_{lp}$ w/ $\mathcal{L}_{t}$ 0.0368 0.0167 8.20 21.29
ST w/ DSR w/ large $\mathcal{L}_{lp}$ w/ $\mathcal{L}_{t}$ 0.2206 0.2082 3.63 31.60
Ours Full 0.0368 0.0174 7.99 10.66
Ablation on Vertex Number for $\mathcal{L}_{lp}$
4,000 (Ours Full) 0.0368 0.0174 7.99 10.66
@@
We conduct extensive experiments to evaluate the significance of each proposed component, and the results are shown in the first part of Tab. 2. The baseline follows R2ET [20], which applies an SDF loss for training but retargets motion frame-by-frame without considering temporal dependencies. Introducing the spatio-temporal (ST) model reduces curvature significantly, demonstrating the effectiveness of temporal modeling in reducing motion jitter. However, the penetration rate increases slightly. Adding the limb penetration constraint ($\mathcal{L}_{lp}$) effectively mitigates interpenetration, reducing the penetration rate from $9.45\%$ to $8.20\%$. Further incorporating the basic smoothness loss ($\mathcal{L}_{t}$) improves temporal consistency without affecting penetration.
Next, we analyze the impact of dense shape representation (DSR). Adding DSR alongside $\mathcal{L}_{lp}$ and $\mathcal{L}_{t}$ further reduces curvature (from 35.74 to 21.29), indicating that shape-aware features help the model predict smoother motion. However, an excessive penalty on penetration, such as applying a large $\mathcal{L}_{lp}$, reduces penetration significantly (to $3.63\%$) but leads to poor motion generalization, as reflected by the sharp increase in MSE and local MSE. Our full model, which balances spatial, temporal, and geometric constraints, achieves a pretty low penetration rate ($7.99\%$) while maintaining low curvature (10.66), demonstrating its overall effectiveness.
In the second part of Tab. 2, we examine the effect of varying the number of sampled vertices for the limb penetration constraint. Reducing the number of sampled vertices from 4,000 (full model) to 500 slightly increases penetration and curvature, but the differences remain minimal, suggesting that our method is robust to different sampling densities. Considering the balance between computational efficiency and accuracy, we choose 4,000 vertices in our final model.
We also conduct the ablation experiment for joint orientation loss, and please refer to Sec. E in the supplemental materials.
@@ 5 Conclusion
In this paper, we aim to achieve three key objectives in motion retargeting: semantics preservation, geometric plausibility, and temporal consistency. We propose STaR, a novel spatio-temporal skinned motion retargeting framework, embedded with a limb penetration constraint and a temporal consistency constraint. These constraints curve the motion manifold for characters of different shapes, while our spatio-temporal model design helps the optimizer find better parameters for this task. Experimental results demonstrate that our STaR achieves a superior balance across all three objectives compared to previous methods.
Limitations and Future Work. LBS is a major factor contributing to interpenetration, and resolving the skinning issue would significantly reduce the penetration rate, as the human body is non-rigid and some penetrations would naturally be replaced by skin deformation. Our current semantics preservation method relies on sparse supervision, and we plan to explore denser and more reliable semantic constraints in future work. Another challenge is missing contact, which we do not address in this paper due to the prevalence of false contact signals—certain motions that do not involve contact may still generate additional contact signals on larger characters. We aim to tackle these challenges in future research.
@@
  • Retargeting motion from real human.
@@
  • Ablation study on global motion.
@@ B Evaluation Metrics
[…]
@@ D User Study
Semantics Preservation $11.43\%$ $8.57\%$ $2.86\%$ $77.14\%$
Geometry Correctness $12.38\%$ $9.52\%$ $4.76\%$ $73.33\%$
Motion Coherence $20.00\%$ $4.76\%$ $4.76\%$ $70.48\%$
Overall Quality $12.38\%$ $8.57\%$ $2.86\%$ $76.19\%$
@@
To ensure a diverse set of responses, we distributed questionnaires and collected feedback from 21 participants. The results, presented in Tab. 3, show that STaR significantly outperforms all baseline methods across all criteria. Notably, $77.14\%$ of participants favored STaR in terms of motion semantics preservation, and $76.19\%$ preferred our results overall. Our model also received $70.48\%$ approval for motion coherence, demonstrating its ability to produce smooth and temporally stable motion trajectories, while achieving $73.33\%$ in geometric correctness, highlighting its effectiveness in preventing interpenetration and ensuring spatial plausibility.
@@ E Ablation Study on Joint Orientation Loss
F Ablation Study on Global Motion
After analyzing the relationship between global motion and character height in the Mixamo dataset [2], we found no clear correlation. To improve global motion prediction, we designed a small decoder; however, the network quickly overfits to the training set, resulting in poor generalization. As shown in Tab. 4, its performance fails to surpass the baseline, which simply normalizes and de-normalizes global motion based on character height.
Table 4: Ablation study on global motion
@@ G Single-pass Motion Retargeting and Separate Motion Retargeting
As shown in Tab. 5, our well-designed spatio-temporal model allows motion retargeting of varying lengths in a single forward pass. The results show no significant difference between separate retargeting and single-pass retargeting.
@@
H Demo Videos

### v2

@@ 4.3 Ablation Study
Table 2: Ablation Study. Baseline is a model which retargets motion frame-by-frame, training with $\mathcal{L}_{sdf}$ from [20]. ST is the spatio-temporal model, and DSR stands for dense shape representation. $\mathcal{L}_{lp}$ is the limb penetration constraint, and $\mathcal{L}_{tc}$ is the temporal consistency constraint. $\mathcal{L}_{t}$ is a basic motion smoothness loss. ✓✓stands for a higher weight for $\mathcal{L}_{lp}$. In the second part of the table, we sample different numbers of vertices for the limb penetration constraint.
@@
[…]
$\mathcal{L}_{sdf}$ 0.0479 0.0321 8.74 66.35
$\mathcal{L}_{sdf}$ 0.0371 0.0179 9.45 7.87
0.0361 0.0167 8.20 35.74
$\mathcal{L}_{t}$ 0.0361 0.0167 8.40 35.74
$\mathcal{L}_{t}$ 0.0368 0.0167 8.20 21.29
✓✓ $\mathcal{L}_{t}$ 0.2206 0.2082 3.63 31.60
0.0368 0.0174 7.99 10.66
Ablation study on vertex number for $\mathcal{L}_{lp}$
4,000 (Ours) 0.0368 0.0174 7.99 10.66
@@
[…]
In the second part of Tab. 2, we examine the effect of varying the number of sampled vertices for the limb penetration constraint. Reducing the number of sampled vertices from 4,000 (ours) to 500 slightly increases penetration and curvature, but the differences remain minimal, suggesting that our method is robust to different sampling densities. Considering the balance between computational efficiency and accuracy, we choose 4,000 vertices in our final model. We conduct further ablation experiments for other modules, and please refer to the supplemental materials.
@@ 5 Conclusion
In this paper, we aim to achieve three key objectives in motion retargeting: semantics preservation, geometric plausibility, and temporal consistency. We propose STaR, a novel spatio-temporal skinned motion retargeting framework, embedded with a limb penetration constraint and a temporal consistency constraint. These constraints modify the motion manifold for characters of different shapes, while our spatio-temporal model design helps the optimizer find better parameters for this task. Experimental results demonstrate that our STaR achieves a superior balance across all three objectives compared to previous methods.
Limitations and Future Work. LBS is a major contributor to interpenetration. Resolving skinning artifacts could significantly lower penetration rates, since some issues stem from rigid-body assumptions that real skin deformation would naturally avoid. We also do not tackle missing contact in this paper, due to frequent false positives. Addressing these limitations will be a focus of our future research.
Acknowledgements
GS acknowledges support by Engineering and Physical Sciences Research Council [grant number EP/Y009800/1], through Keystone project funding from Responsible AI UK (KP0016). This research utilized Queen Mary’s Andrena HPC facility, supported by QMUL Research-IT.
@@ Acknowledgements
  • Real-world applications: retargeting motion from real humans.
@@
  • Ablation study on global motion prediction.
@@
  • Ablation study on Dense Shape Representation.
  • Efficiency of limb penetration constraint module.
@@ B Evaluation Metrics
$\mathbf{Pen\ Rate}=\frac{\text{Number of penetrated limb vertices}}{\text{Number of all limb vertices}}\times 100\%.$ (9)
@@ D User Study
[…]
@@
[…]
@@ E Ablation Study on Joint Orientation Loss
F Ablation Study on Global Motion Prediction
After examining the relationship between global motion and character height in the Mixamo dataset [2], we observed no statistically significant correlation. After introducing a compact decoder to enhance global motion prediction, we observed rapid overfitting to the training set, resulting in poor generalization. As shown in Tab. 4, its performance does not exceed the baseline, which simply normalizes and denormalizes global motion relative to character height.
Table 4: Ablation study on global motion prediction. Since the global motion prediction module will only affect the MSE metric, the other three metrics are omitted for clarity. Please note that this study evaluates a preliminary variant—not our final model configuration.
@@ G Single-pass Motion Retargeting and Separate Motion Retargeting
As presented in Tab. 5, our well-designed spatio-temporal model supports motion retargeting of varying sequence lengths within a single forward pass. The results show no significant difference between separate retargeting and single-pass retargeting.
@@
H Ablation study on Dense Shape Representation
For shape representation, we exclude skeleton data and skeleton-based shape information due to their limited utility and potential drawbacks. The skeleton bounding box’s dimensions [20] provide limited shape details, which are insufficient for effective motion retargeting that minimizes penetration. This limitation arises because motion is influenced by the shapes of limbs and other body parts. We show 3 results as proof in Tab. 6: (1) skeleton data only; (2) our DSR; (3) their combination. Comparing (1) and (2), skeleton data is inadequate for preventing penetration. Additionally, (3) shows that integrating skeleton data with point clouds does not enhance results.
In DSR, the point cloud is NOT separated before being fed into the point cloud transformer [7], and the $K$-channel shape representation is derived via a compact MLP from the comprehensive geometric feature produced by the transformer. Unlike frame-by-frame methods [20, 19] that rely on spatially aligned skeleton data, our approach extracts both global and local information for each joint. This enables the spatial and temporal transformers to access comprehensive shape information, particularly benefiting the temporal network. In Tab. 6 (4), we show the result of extracting a separate point cloud for each joint as proof. Comparing (2) and (4), separating the points does not benefit the retargeting pipeline.
Table 6: Ablation study on Dense Shape Representation.
[…]
Global Motion Prediction
(1) Skeleton Info 0.0367 0.0175 9.42 10.11
(2) DSR (Ours) 0.0368 0.0174 7.99 10.66
(3) Skeleton Info + DSR 0.0366 0.0174 8.22 10.05
(4) Separate Point Cloud 0.0934 0.0783 13.45 8.70
I Efficiency of the Limb Penetration Constraint Module
In Tab. 7, we compare the training speeds of (1) the SDF loss $\mathcal{L}_{sdf}$ from R2ET [20], (2) modified SDF loss [20] including limb-limb penetration, and (3) our limb penetration constraint $\mathcal{L}_{lp}$. Using the same number of vertices, we measured the time per iteration on an RTX 4090 graphics card. Our limb penetration loss is approximately 8 times faster compared with the SDF loss from [20].
Table 7: Loss Efficiency.
[…]
$\mathcal{L}_{sdf}$ [20] 127.66
$\mathcal{L}_{sdf}$ [20] w/ limb 133.77
$\mathcal{L}_{lp}$ (Ours) 16.32
J Demo Videos